What is The electric field E (peak value) in a microwave beam density 100W/m² travelling in free space
1
Expert's answer
2019-04-16T09:43:06-0400

The wave intensity is


"I=c \epsilon_0 E_m^2"

Thus,


"E_m=\sqrt{\frac{I}{c \epsilon_0 }}=\sqrt{\frac{100}{3\cdot 10^8 (8.85\cdot 10^{-12} ) }}=194 \frac{V}{m}"